write a quadratic equation in standard form having the given solutions 4i,-4i
Answer by ankor@dixie-net.com(22740)

write a quadratic equation in standard form having the given solutions 4i,-4i
x = +/-4i
x^2 = 16(i^2)
x^2 = 16(-1)
x^2 = -16
x^2 + 16 = 0
y = x^2 + 16; is the standard form